OpenStudy (anonymous):

[6.05] Your piggy bank has a total of 40 coins in it; some are dimes and some are nickels. If you have a total of $3.50, how many dimes do you have? 10 20 30 40

OpenStudy (anonymous):

D=dimes N=Nickels D+N=40 (This is for the number of coins) D*10 + N*5 = 3.50 Dimes cost 10 cents that's why i multiply 10 and Nickles cost 5 cents that's why I multiply by 5. Add them together and we have 3.50 Now solve the system of equations.

OpenStudy (anonymous):

is the answer 10 dimes?

OpenStudy (anonymous):

I don't know did you get 10 dimes when you solved it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What's 10*0.10=?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

ok now you have 30 nickles left what's 30* 0.05=?

OpenStudy (anonymous):

1.5

OpenStudy (anonymous):

1+1.5=?

OpenStudy (anonymous):

2.5

OpenStudy (anonymous):

so 10 dimes is not the answer.

OpenStudy (anonymous):

Try solving the system of equations again D+N=40 (This is for the number of coins) D*0.10 + N*0.05 = 3.50

OpenStudy (anonymous):

do i just add 0.10 + 0.05?

OpenStudy (anonymous):

Isolate N for both equation and then substitute it to get rid of N and you are left with D and then you solve for D For example: D+N=40 --> N=40-D D*0.10 + N*0.05 = 3.50 --> N = (3.50 - D*0.10) / 0.05 Now substitute: 40-D = (3.50 - D*0.10) / 0.05 Now solve for D

OpenStudy (anonymous):

40-D=3.50-Dx0.10/0.05 -3.50 -3.50 D=36.5-Dx0.10/0.05 D=36.5-Dx0.10/0.05 D=0.10+730 D+730.1

OpenStudy (anonymous):

Thanks Romero